OpenStudy (anonymous):

7. Which ordered pair is a solution of the equation y = –9x + 4? (1 point) (10, –86) (–4, –58) (6, –41) (–6, 57)

OpenStudy (nurali):

y = -9x + 4 The first is (10 -86), so try x = 10 and see whether y turns out to be -86. y = -9(10) + 4 = -90 + 4 = -86 Yes it does, so (10, -86) is a solution.